Let X1 and X2 be uniformaly distributed over the interval zero to one, (a) What is the joint probability density function of Y1 = X1+X2 and Y2 = X2-X1? (b) Are Y1 and Y2 independent? Explain.
Answer #1

Given X and Y are uniformly distributed between 0 and 1, so

f(3) = 1:0 <I< 1,0; otherwise.if we consider X1 and Y1 jointly uni- formly distributed thanf(11,12) = 1; 0 < (11,12) <10; oth

So we use Jacobian to find the PDF for  (Y,Y) from the PDF of (x1, X₂) . The solution is as followslet y = x1 + x2 and Y2 2 X2-XI, ow, these two funtions are one-to-one and continuons and denivative (pantial) enest and alsoang 592 9,-4,22 0) ( 9+4 = 2 ,- - 2 h (a,02) 9 ₂ (ag, 2) 242 hi (4, 9) = x1 halb,,Y₂) = x2 fly, I₂) = f (h. 19, 2), 5 (81, 8)fy, (3,) = 1 1 (4) dy Oy ocy, ci y + 2 V (1/2) dyz, 849,22 9,2 ocycl { [921 of 5,4 1 - +2 lay, L2 2192), -2 2 x+24) ocy, ci c2-42 04 3 day, occi +7, (3) Ocy Y2 2+42 LY 20 2-42 OLY₂ 21 - 3125, 12, ocyan yz 2742 - 24,20 { [y]2492 -129240 02 - 12-92-4]h(y1, y2) = 5 + fy(y)fY,(92),so Y and Y2 are not independent, if they were independent then it should satisfy h(y1, y2) = fy,